# Axis Tour: Word Tour Determines the Order of Axes in ICA-transformed Embeddings

### v2

@@ 1 Introduction
Figure 1: Scatterplots of normalized ICA-transformed word embeddings whose axes are ordered by Axis Tour and Skewness Sort. In the upper part, Axis Tour is applied to 300-dimensional GloVe, with nine consecutive axes arranged counterclockwise. In the lower part, these nine axes are arranged clockwise in descending order of skewness. The embeddings are projected onto two dimensions along these axes. The top five embeddings on each axis are labeled by their words. Each word is assigned the color of the axis with the highest value. In both cases, words that cross the horizontal axes are not displayed. Refer to Appendix A for more details.
Embedding is an important tool in natural language processing, but interpreting high-dimensional embeddings is challenging. To address this, Independent Component Analysis (ICA) Hyvärinen and Oja (2000) is an effective solution Mareček et al. (2020); Musil and Mareček (2022); Yamagiwa et al. (2023). ICA-transformed embeddings reveal interpretable semantic axes; however, the order of these axes are arbitrary Hyvärinen et al. (2001b). In this study, inspired by a one-dimensional word embeddings, Word Tour Sato (2022), which leverages the Traveling Salesman Problem (TSP), we aim to improve the clarity of the word embedding space by maximizing the semantic continuity of the axes.
Fig. 1 shows two sets of two-dimensional projections of word embeddings: one is ordered by Axis Tour and the other is sorted in descending order of skewness (Skewness Sort). In Axis Tour, the top words of the axes are far from the center, with the meanings of the axes changing continuously. Conversely, in Skewness Sort, the top words are nearer to the center, and the axes with different meanings are placed adjacently. In fact, the average distance from the origin to the top words in Fig. 1 is $0.76$ in Axis Tour, as opposed to $0.61$ in Skewness Sort.
We also assume that the consecutive axes in the Axis Tour embeddings can be considered as a subspace whose axes have similar meanings. Based on this, we project each subspace onto a single dimension for dimensionality reduction. We show through experiments that Axis Tour yields better or comparable low-dimensional embeddings compared to both PCA and ICA.
@@ 2 Related work
Relevant for our study is Topographic ICA (TICA) Kohonen (2001); Hyvärinen et al. (2001a). TICA relaxes the assumption of statistical independence and assumes higher-order correlations between adjacent axes. TICA then estimates the order of the axes. Unlike TICA, Axis Tour applies to ordinary ICA-transformed embeddings and uses the embeddings themselves to measure axis similarity. For more details on TICA, refer to Appendix F.
@@ 3 Background
The pre-trained word embedding matrix is given by $\mathbf{X}=[\mathbf{x}_{1},\cdots,\mathbf{x}_{n}]^{\top}\in\mathbb{R}^{n\times d}$, where $\mathbf{X}$ is centered (i.e., the mean of each column is zero). Here, $\mathbf{x}_{i}\in\mathbb{R}^{d}$ represents the word embedding of the $i$th word.
@@ 3.1 ICA-transformed word embeddings
where $\mathbf{S}$ is whitened (i.e., the variance of the columns is $1$ and their correlation is 00). The columns of $\mathbf{S}$ are called independent components [^1] Flip axis signs as needed to make the skewness positive.. While $\mathbf{S}$ has interpretable semantic axes Mareček et al. (2020); Musil and Mareček (2022); Yamagiwa et al. (2023), the order of these axes are arbitrary Hyvärinen et al. (2001b).
@@ 3.2 Word Tour
Let $\mathcal{P}([n])$ be the set of all permutations of $[n]$, where $[n]=\{1,\cdots,n\}$. Word Tour Sato (2022) is a one-dimensional word embeddings method which solves the following Traveling Salesman Problem (TSP):
@@
The resulting one-dimensional embeddings have similar meanings if they are close in order.
@@ 4 Axis Tour
This section explains Axis Tour and the dimensionality reduction method using Axis Tour. As mentioned in Section 3.2, $[d]=\{1,\cdots,d\}$ and $\mathcal{P}([d])$ is the set of all permutations of $[d]$.
@@ 4.1 Definition of axis embedding
In preparation, we define the matrix $\hat{\mathbf{S}}\in\mathbb{R}^{n\times d}$ as the normalization of the embeddings $\mathbf{S}$ with row vectors $\hat{\mathbf{s}}_{i}=\mathbf{s}_{i}/\|\mathbf{s}_{i}\|$. Here, the $i$th word embedding of $\mathbf{S}$ and $\hat{\mathbf{S}}$ are denoted by $\mathbf{s}_{i},\hat{\mathbf{s}}_{i}\in\mathbb{R}^{d}$, respectively. We compare the elements of the $\ell$th axis of $\hat{\mathbf{S}}$ and denote the index set of words corresponding to the top $k$ elements as $\text{Top}_{k}^{\ell}$. We then define the $\ell$th axis embedding $\mathbf{v}_{\ell}$ for $\mathbf{S}$ as follows:
@@
As we saw in Fig. 1, since the meaning of an axis can be interpreted from the top words, $\mathbf{v}_{\ell}$ can be considered as the embedding that represents the meaning of the $\ell$th axis of $\mathbf{S}$.
4.2 Determining the order of axes
Axis Tour is a method that uses $\mathbf{v}_{\ell}$ in (3) to perform Word Tour and determines the order of axes in ICA-transformed word embeddings. In Axis Tour, the cost between the axis embeddings $\mathbf{v}_{\ell},\mathbf{v}_{m}$ used for TSP is defined by $1-\cos{(\mathbf{v}_{\ell},\mathbf{v}_{m})}$ instead of $\|\mathbf{v}_{\ell}-\mathbf{v}_{m}\|$. This approach then maximizes the sum of cosine similarities between adjacent axis embeddings. Therefore, the problem is formulated as follows [^2] Note that due to the cyclic nature of $\tau$, we ensure that $\cos(\mathbf{v}_{\tau_{1}},\mathbf{v}_{\tau_{d}})$ is the smallest of the cosine similarities.:
4.3 Dimensionality reduction
[…]
First, we consider dimensionality reduction for $I_{r}$. To do this, we define a unit vector $\mathbf{f}_{r}:=(f_{r}^{(\ell)})_{\ell=1}^{d}\in\mathbb{R}_{\geq 0}^{d}$ for each $I_{r}$ as follows:
Next, we define the matrix $\mathbf{F}:=[\mathbf{f}_{1},\cdots,\mathbf{f}_{p}]\in\mathbb{R}^{d\times p}$. Then $\mathbf{T}\mathbf{F}\in\mathbb{R}^{n\times p}$ represents the concatenated projections, that is, a dimensionality reduction of the $d$-dimensional embeddings $\mathbf{T}$ to $p$ dimensions. For more details, refer to Appendix B.
5 Experiments
@@
Table 1: Semantic continuity of axes by Axis Tour for normalized ICA-transformed embeddings. We apply Axis Tour to 300-dimensional GloVe and show the top five words for each axis. See Appendix E.1 for all axis results.
@@
Figure 2: Histogram of $\cos(\mathbf{v}_{\ell},\mathbf{v}_{\ell+1})$. As an additional baseline, we sampled 300 random words from the Random Order embeddings, ordered them, and showed the histogram. The dashed lines are the average similarity for each method. The distribution for Axis Tour shifts towards a more positive mean. Conversely, others roughly follow the normal distribution with means that are close to 0. For more details, refer to Appendix C.
@@
Figure 3: Relationship between the skewness and the average of consecutive two cosines (i.e., $(\cos(\mathbf{v}_{\ell-1},\mathbf{v}_{\ell})+\cos(\mathbf{v}_{\ell},\mathbf{v}_{\ell+1}))/2$) for (a) Axis Tour and (b) Skewness Sort. The left shows the skewness and the average of two cosines for both $y$-axes, and the right shows the scatter plot of these values. Spearman’s rank correlation is 0.43 for Axis Tour, while it is 0.04 for Skewness Sort.
Similar to the Word Tour experiments, we used 300-dimensional GloVe  Pennington et al. (2014) with $n=400{,}000$, and the LKH solver [^5] The LKH solver is an implementation of Lin-Kernighan algorithm Lin and Kernighan (1973); Helsgaun (2000). Helsgaun (2018) as the TSP solver.
For ICA, we used FastICA Hyvärinen (1999) from scikit-learn Pedregosa et al. (2011), setting iterations to $10{,}000$ and tolerance to $10^{-10}$, consistent with Yamagiwa et al. (2023). We computed the axis embeddings [^6] See Appendix E.3 for a discussion of the choice of $k$. $\mathbf{v}_{\ell}$ in (3) with $k=100$. For baselines, we used whitened PCA-transformed embeddings [^7] Whitened ICA-transformed embeddings are obtained by applying an orthogonal matrix to these embeddings., along with two types of whitened ICA-transformed embeddings [^8] See Appendix F for comparisons of Axis Tour and TICA.: Random Order, which randomly flips the sign of the axes in $\mathbf{S}$ and randomizes the order of the axes, and Skewness Sort, which sorts the axes of $\mathbf{S}$ in descending order of skewness. See Appendix E for additional experiments, including those of other embeddings.
5.1 Qualitative observation
Table 1 shows the illustrative examples of the meaning of countries, law, and art by the consecutive axes of the Axis Tour embeddings. These examples illustrate how the meanings of the axes change continuously. For example, in the top row, the axis meaning changes from Eastern Europe to Germany and France, followed by Canada (which shares a connection with France), then to Australia (English-speaking regions), the regions in England, and finally to soccer (a popular sport in England), demonstrating geographic and cultural continuity.
5.2 Cosine similarity between adjacent axis embeddings
Fig. 2 shows the histograms of $\cos(\mathbf{v}_{\ell},\mathbf{v}_{\ell+1})$ for Axis Tour and the baselines. In Axis Tour, the values of $\cos(\mathbf{v}_{\ell},\mathbf{v}_{\ell+1})$ are consistently higher, while this trend is not observed in the other baselines. This result is consistent with the formulation in (4).
For Axis Tour and Skewness Sort, Fig. 3 illustrates the relationship between the skewness and the average of consecutive two cosine similarities [^9] Cosine similarity is not defined for the $i$th axis itself.. In Fig. 3(a) for Axis Tour, it is interesting to note that these two values are correlated. However, no such trend is observed in Fig. 3(b) for Skewness Sort. This indicates that in Axis Tour, the axes with higher skewness are more likely to be ordered adjacent to the axes with higher similarity.
@@ 5.3 Dimensionality reduction: analogy, word similarity, and categorization tasks
Using Word Embedding Benchmark Jastrzebski et al. (2017) [^10] https://github.com/kudkudak/word-embeddings-benchmarks., we evaluated the performance of dimensionality reduction in analogy, word similarity, and categorization tasks. PCA selects the axes in descending order of eigenvalue. Random Order and Skewness Sort select the axes sequentially from the first to last. Axis Tour adopts the dimensionality reduction [^11] Fig. 7 in Appendix E.2 shows the results for different $\alpha$. in Section 4.3 with $\alpha=1/3$.
Fig. 4 shows that the dimensionality reduction for Axis Tour is better than or comparable to the baselines for most dimensions in each task. This result suggests that the dimensionality reduction for Axis Tour efficiently merges the axes with similar meanings. For more details, refer to Appendix D.

### v3

@@ 1 Introduction
Figure 1: Scatterplots of normalized ICA-transformed word embeddings whose axes are ordered by Axis Tour and Skewness Sort. In the upper part, Axis Tour is applied to 300-dimensional GloVe, with nine consecutive axes arranged counterclockwise. In the lower part, these nine axes are rearranged clockwise in descending order of skewness. The embeddings are projected onto two dimensions along these axes. The top five embeddings on each axis are labeled by their words. Each word is assigned the color of the axis with the highest value. In both cases, words that cross the horizontal axes are not displayed. Refer to Appendix A for more details.
Embedding is an important tool in natural language processing, but interpreting high-dimensional embeddings is challenging. To address this, Independent Component Analysis (ICA) Hyvärinen and Oja (2000) offers an effective solution Mareček et al. (2020); Musil and Mareček (2024); Yamagiwa et al. (2023). ICA-transformed embeddings reveal interpretable semantic axes; however, the order of these axes is arbitrary Hyvärinen et al. (2001b). In this study, inspired by a one-dimensional word embedding method, Word Tour Sato (2022), which leverages the Traveling Salesman Problem (TSP), we aim to improve the clarity of the word embedding space by maximizing the semantic continuity of the axes.
Figure 1 shows two sets of two-dimensional projections of word embeddings: one is ordered by Axis Tour (our proposal), and the other is sorted in descending order of skewness (Skewness Sort). In Axis Tour, the top words of the axes are positioned farther from the center, with the meanings of the axes changing continuously. Conversely, in Skewness Sort, the top words are closer to the center, and the axes with different meanings are placed adjacently. In fact, the average distance from the origin to the top words in Fig. 1 is $0.76$ in Axis Tour, compared to $0.61$ in Skewness Sort.
We also assume that the consecutive axes in the Axis Tour embeddings can be considered a subspace with axes that have similar meanings. Based on this idea, we project each subspace onto a single dimension for dimensionality reduction. We show through experiments that Axis Tour yields better or comparable low-dimensional embeddings compared to both PCA and ICA.
@@ 2 Related work
Research on interpreting embeddings by focusing on axes representing opposing concepts (e.g., cold vs. hot, soft vs. hard) is also actively pursued. Approaches such as SemAxis An et al. (2018), POLAR Mathew et al. (2020), and FrameAxis Kwak et al. (2021) deal with static embeddings, while BiImp Senel et al. (2022) and SensePOLAR Engler et al. (2022) deal with dynamic embeddings. In particular, Section 5.3 provides a comparison of the Axis Tour embeddings and those from POLAR.
Relevant to our study is Topographic ICA (TICA) Kohonen (2001); Hyvärinen et al. (2001a). TICA relaxes the assumption of statistical independence and assumes higher-order correlations between adjacent axes. then estimates the order of the axes. Unlike TICA, Axis Tour is applied to ordinary ICA-transformed embeddings and uses the embeddings themselves to measure axis similarity. For more details on TICA, refer to Appendix F.
@@ 3 Background
The pre-trained word embedding matrix is given by $\mathbf{X}=[\mathbf{x}_{1},\ldots,\mathbf{x}_{n}]^{\top}\in\mathbb{R}^{n\times d}$, where $\mathbf{X}$ is centered (i.e., the mean of each column is zero). Here, $\mathbf{x}_{i}\in\mathbb{R}^{d}$ represents the word embedding of the $i$-th word.
@@ 3.1 ICA-transformed word embeddings
where $\mathbf{S}$ is whitened (i.e., the variances of the columns are $1$ and their correlations are all 00). The columns of $\mathbf{S}$ are called independent components [^1] Unless otherwise noted, flip the sign of each axis as needed so that the skewness is positive.. While $\mathbf{S}$ has interpretable semantic axes Mareček et al. (2020); Musil and Mareček (2024); Yamagiwa et al. (2023), the order of these axes are arbitrary Hyvärinen et al. (2001b).
@@ 3.2 Word Tour
Let $\mathcal{P}([n])$ be the set of all permutations of $[n]$, where $[n]=\{1,\ldots,n\}$. Word Tour Sato (2022) is a one-dimensional word embedding method that solves the following Traveling Salesman Problem (TSP):
@@
The resulting one-dimensional embeddings have similar meanings when they are close in order.
@@ 4 Axis Tour
This section explains Axis Tour and the dimensionality reduction method using Axis Tour. As mentioned in Section 3.2, $[d]=\{1,\ldots,d\}$ and $\mathcal{P}([d])$ is the set of all permutations of $[d]$.
@@ 4.1 Definition of axis embedding
In preparation, we define the normalized ICA-transformed embeddings $\hat{\mathbf{S}}\in\mathbb{R}^{n\times d}$ as the normalization of the embeddings $\mathbf{S}$, where the row vectors are given by $\hat{\mathbf{s}}_{i}=\mathbf{s}_{i}/\|\mathbf{s}_{i}\|$. Here, the $i$-th word embeddings of $\mathbf{S}$ and $\hat{\mathbf{S}}$ are denoted by $\mathbf{s}_{i}$ and $\hat{\mathbf{s}}_{i}\in\mathbb{R}^{d}$, respectively. We compare the elements of the $\ell$-th axis of $\hat{\mathbf{S}}$ and denote the index set of words corresponding to the top $k$ elements as $\text{Top}_{k}^{\ell}$. We then define the $\ell$-th axis embedding $\mathbf{v}_{\ell}$ for $\mathbf{S}$ as follows:
@@
As we saw in Fig. 1, since the meaning of an axis can be interpreted from the top words, $\mathbf{v}_{\ell}$ can be considered the embedding that represents the meaning of the $\ell$-th axis of $\mathbf{S}$.
@@
Table 1: Semantic continuity of axes by Axis Tour for normalized ICA-transformed embeddings. We apply Axis Tour to 300-dimensional GloVe and show the top five words for each axis. See Appendix E.1 for all axes results.
4.2 Determining the order of axes
Axis Tour is a method that uses $\mathbf{v}_{\ell}$ in (3) to perform Word Tour and determines the order of axes in ICA-transformed word embeddings. In Axis Tour, the cost between the axis embeddings $\mathbf{v}_{\ell}$ and $\mathbf{v}_{m}$ for the TSP is defined by $1-\cos{(\mathbf{v}_{\ell},\mathbf{v}_{m})}$ instead of $\|\mathbf{v}_{\ell}-\mathbf{v}_{m}\|$. This approach then maximizes the sum of cosine similarities between adjacent axis embeddings. Therefore, the problem is formulated as follows [^2] Note that due to the cyclic nature of $\tau$, we set $\tau_{1}$ such that $\cos(\mathbf{v}_{\tau_{1}},\mathbf{v}_{\tau_{d}})$ is the smallest of the cosine similarities.:
4.3 Dimensionality reduction
[…]
First, we consider reducing the dimensionality of $\mathbf{T}$ along $I_{r}$, $r=1,\ldots,p$. To do this, we define a unit vector $\mathbf{f}_{r}:=(f_{r}^{(\ell)})_{\ell=1}^{d}\in\mathbb{R}_{\geq 0}^{d}$ for each $I_{r}$ as follows:
Next, we define the matrix $\mathbf{F}:=[\mathbf{f}_{1},\ldots,\mathbf{f}_{p}]\in\mathbb{R}^{d\times p}$. Then $\mathbf{T}\mathbf{F}\in\mathbb{R}^{n\times p}$ represents the concatenated projections, that is, a dimensionality reduction of the $d$-dimensional embeddings $\mathbf{T}$ to $p$ dimensions. For more details, refer to Appendix B.
5 Experiments
Similar to the Word Tour experiments, we used 300-dimensional GloVe  Pennington et al. (2014) with $n=400{,}000$, and the LKH solver [^5] The LKH solver is an implementation of Lin-Kernighan algorithm Lin and Kernighan (1973); Helsgaun (2000). Helsgaun (2018) for the optimization of the TSP.
For ICA, we used FastICA Hyvärinen (1999) from scikit-learn Pedregosa et al. (2011), setting the iterations to $10{,}000$ and the tolerance to $10^{-10}$, consistent with Yamagiwa et al. (2023). We computed the axis embeddings [^6] See Appendix E.3 for a discussion of the choice of $k$. $\mathbf{v}_{\ell}$ in (3) with $k=100$. For baselines, we used whitened PCA-transformed embeddings [^7] Whitened ICA-transformed embeddings are obtained by applying an orthogonal matrix to these embeddings., along with two types of whitened ICA-transformed embeddings [^8] See Appendix F for comparisons of Axis Tour and TICA.: Random Order, which randomly flips the sign of the axes in $\mathbf{S}$ and randomizes the order of the axes, and Skewness Sort, which sorts the axes of $\mathbf{S}$ in descending order of skewness. See Appendix E for additional experiments, including those of other embeddings.
5.1 Qualitative observation of semantic continuity in axis order
Table 1 presents an illustrative example of consecutive axes of the Axis Tour embeddings, where the three rows correspond to the meanings of countries, law, and art, respectively. We observe that the meanings of the axes change continuously. For instance, in the top row, the axis meaning shifts from Eastern Europe to Germany and France, followed by Canada (which shares a connection with France), then to Australia (English-speaking regions), the regions in England, and finally to soccer (a popular sport in England), demonstrating geographic and cultural continuity.
5.2 Quantitative evaluation of semantic continuity in axis order
This section quantitatively evaluates the semantic continuity of the axes in Axis Tour embeddings.
5.2.1 Evaluation by cosine similarity
First, we evaluate the two orderings of axes shown in Fig. 1. The semantic continuity of these axes is assessed by calculating the average cosine similarity between adjacent axis embeddings. For Axis Tour, the average cosine similarity is 0.269, but it decreases to 0.185 when these axes are rearranged by skewness, confirming the higher semantic continuity of the axes in Axis Tour.
Next, we consider the whole $d\,(=300)$ consecutive axes. Figure 2 shows the histograms of $\cos(\mathbf{v}_{\ell},\mathbf{v}_{\ell+1})$ for Axis Tour and the baselines. In Axis Tour, the values of $\cos(\mathbf{v}_{\ell},\mathbf{v}_{\ell+1})$ are consistently higher, while this trend is not observed in the other baselines. The average cosine similarity is 0.244 for Axis Tour, while it is only 0.017 for Skewness Sort. This result is consistent with the formulation in (4).
5.2.2 Evaluation by GPT models
We also evaluate the semantic continuity of axes for Axis Tour and Skewness Sort using the OpenAI API. By focusing on the common axes in each embedding, we ask the model to determine, based on the top 10 words, whether the next axis of Axis Tour or that of Skewness Sort is more semantically related. The number of queries corresponds to $d\,(=300)$, and we use four GPT models: GPT-3.5 Turbo, GPT-4 Turbo, GPT-4o, and GPT-4o mini (see Appendix G for model versions and prompts).
As shown in Fig. 3, Axis Tour has a greater number of related axes compared to Skewness Sort for each model, implying more continuous changes in axis meanings. The smallest difference was observed with GPT-3.5 Turbo, the least performant model. For the other models (i.e., the GPT-4 models), the difference was at least four times larger.
@@ 5.2.2 Evaluation by GPT models
Figure 2: Histogram of $\cos(\mathbf{v}_{\ell},\mathbf{v}_{\ell+1})$. As an additional baseline, we sampled 300 random words from the Random Order embeddings and arranged them in random order. The dashed lines represent the average similarity for each method. The distribution for Axis Tour shifts towards a more positive mean, while the others roughly follow a normal distribution with means close to 0. For more details, refer to Appendix C.
Figure 3: Comparison of the number of related axes in the GPT models. In each model, Axis Tour exhibits a greater number of related axes than Skewness Sort.
@@
Figure 5: Relationship between the skewness $\gamma_{\ell}$ and the average of two consecutive cosines $(\cos(\mathbf{v}_{\ell-1},\mathbf{v}_{\ell})+\cos(\mathbf{v}_{\ell},\mathbf{v}_{\ell+1}))/2$ for all the axes $\ell=1,\ldots,d$ in (a) Axis Tour and (b) Skewness Sort. The left plot shows the skewness and the average of two cosines on both $y$-axes, while the right plot shows the scatter plot of these values. Spearman’s rank correlation is 0.43 for Axis Tour, while it is 0.04 for Skewness Sort.
@@ 5.3 Dimensionality reduction: analogy, word similarity, and categorization tasks
Using Word Embedding Benchmark Jastrzebski et al. (2017) [^9] https://github.com/kudkudak/word-embeddings-benchmarks., we evaluate the performance of dimensionality reduction in analogy, word similarity, and categorization tasks. PCA selects the axes in descending order of eigenvalue. Random Order and Skewness Sort select the axes sequentially from first to last. Axis Tour adopts the dimensionality reduction [^10] Fig. 8 in Appendix E.2 shows the results for different $\alpha$. in Section 4.3 with $\alpha=1/3$.
We consider the original GloVe embeddings $\mathbf{X}$ as well as the embeddings obtained by applying POLAR to $\mathbf{X}$. POLAR is a method that uses pairs of words with opposite meanings and finds the axes where these words are positioned at the opposite ends. There are three methods for selecting the axes: Random Selection, Variance Maximization, and Orthogonality Maximization. We used the publicly available code for our experiments. For both the original and POLAR-applied embeddings, axes were selected sequentially from first to last.
Figure 4 shows that the dimensionality reduction with the ordering in Axis Tour is better than or comparable to the baselines for most dimensionalities in each task. This result suggests that Axis Tour efficiently merges axes with similar meanings. For more details, refer to Appendix D.
6 Discussion
We confirmed through both qualitative and quantitative experiments that the axis order determined by Axis Tour exhibits high semantic continuity, and its effectiveness was also validated in the downstream task of dimensionality reduction.
There are two advantages to ordering the axes of ICA-transformed embeddings using Axis Tour. First, as shown in Fig. 1, when projecting the embedding space, the scatterplot becomes easier to interpret and more visually accessible because higher-ranking words on each axis are farther from the origin. Second, although the Axis Tour embeddings $\mathbf{T}$ are the same size as the ICA-transformed embeddings $\mathbf{S}$, the axis order in Axis Tour preserves information about the similarities between the axes.
An interesting relationship between semantic continuity and skewness is shown in Fig. 5, where the semantic continuity of axis $\ell$ is measured by the average of two cosines. In Axis Tour, there is a high correlation between skewness and semantic continuity, while in Skewness Sort, they are almost uncorrelated. A large skewness indicates that an axis has a distinctive meaning, and in Axis Tour, this seems to contribute to semantic continuity.